OpenStudy (anonymous):

Given that , 20=a+b+c+d 336=a×b×c×d b=d+a+d c=a+d+a d+d=a+a+a b+b=c+a+c a+c=d+d+d what is the value of a,b,c, and d?

OpenStudy (dayakar):

a+b+c+d=20 ---------(1) abcd = 336-------------(2) b= a+2d-----------------(3) c= 2a+d --------------(4) 2d =3a ----------------(5) 2b = 2c +a --------------(6) a +c = 3d ------------(7) do u understand this then we can proceed

OpenStudy (anonymous):

What computers are for. Refer to the attachment.

OpenStudy (anonymous):

I understand there